# Combining Evidence Across Filtrations

### v4

@@ D.1 Illustrative example: Sequentially testing a bounded IID mean
For this test martingale, it is possible to derive the conditions with which the running maximum of the test martingale $(M_{t})_{t\geq 0}$ tends to infinity as $t\to\infty$, even under the null. Let $M_{t}^{*}=\sup_{i\leq t}M_{i}$ denote the running maximum of the test martingale. The result below follows from an anti-concentration lemma due to Ramdas et al., (2020).
Proposition 1.
Suppose that the $\mathbb{F}$-predictable sequence $(\lambda_{t})_{t\geq 0}$ are chosen such that, under each $P\in\mathcal{P}^{\mu}$, the sum of quadratic variations in the data, weighted by $\lambda_{t}^{2}$, tends to infinity:
@@
A practically relevant case is when the sequence decays at a rate of $\left({\sqrt{t\log t}}\right)^{-1}$ (or a slower rate). For example, Waudby-Smith and Ramdas, (2024) recommends using $\hat{\lambda}_{t}\propto\left(\sqrt{\hat{\sigma}_{t-1}^{2}t\log(t+1)}\right)^{-1}$, where $\hat{\sigma}_{t-1}^{2}$ is a $\mathcal{F}_{t-1}$-measurable, consistent estimate of the variance $\sigma^{2}$. In such cases, we may use classical generalizations of the SLLN to martingales involving predictable weights (see, e.g., theorem 2.18 of Hall and Heyde,, 1980) to show that condition (41) is met.
@@
We also know, by construction, each $Y_{t}$ is bounded by the constant $2$. Writing $\epsilon=1$, we have
@@
These two conditions, (44) and (45), are sufficient for an “anti-concentration” lemma for test martingales (Ramdas et al.,, 2020, lemma 33). In particular, the conditions imply that
$P\left(\sup_{t}M_{t}\geq x\right)\geq\frac{1}{(1+\epsilon)x}=\frac{1}{2x},\quad\forall x\in(1,\infty).$ (46)
This is enough to show that the running maximum process $M_{t}^{*}=\sup_{i\leq t}M_{i}$ tends to infinity in expectation. More precisely,
$\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}^{*}}\right]\geq\mathbb{E}_{{P}}\left[{\lim_{t\to\infty}M_{t}^{*}}\right]=\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]=\int_{1}^{\infty}P\left(\sup_{t}M_{t}\geq x\right)dx=\int_{1}^{\infty}\frac{dx}{2x}=\infty,$ (47)
@@ D.2 When does this happen for test martingales (and e-processes)?
This phenomenon is not specific to the bounded IID mean testing problem—in fact, we may apply the anti-concentration lemma to any test martingale whose sum of quadratic variations is infinite. Many test martingales used in practice will satisfy this requirement: game-theoretically, as long as one does not decrease one’s bets “too quickly” (or stop betting entirely), which is the case for many powerful test martingales and e-processes, the variance in data will force the sum to infinity. Note that, for test martingales having the form (40), the bets correspond to the predictable sequence $(\lambda_{t})_{t\geq 1}$. More generally for e-processes, we remark that e-processes for composite nulls $\mathcal{P}$ are often constructed as the minimum of powerful test martingales for each null distribution $P\in\mathcal{P}$ (e.g., the UI e-process of Wasserman et al.,, 2020).
@@
By the first half of the anti-concentration lemma (Ramdas et al.,, 2020, lemma 33), we know
@@
Now, suppose for contradiction that $\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]$ were finite. Denoting $M_{\infty}^{*}=\lim_{t\to\infty}M_{t}^{*}=\sup_{t}M_{t}$, we see that $(M_{t})_{t\geq 0}$ is dominated by $M_{\infty}^{*}\in L^{1}$, and thus it is uniformly integrable (see, e.g., theorem 4.6.7 of Durrett,, 2019). This then implies that the martingale $(M_{t})_{t\geq 0}$ converges to $M_{\infty}$ in $L^{1}$. In particular, $\mathbb{E}_{{P}}\left[{M_{\infty}}\right]=\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}}\right]=\mathbb{E}_{{P}}\left[{M_{0}}\right]=1$, which contradicts equation (49). Therefore, $\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]$ must be infinite, and by Fatou’s lemma, $\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}^{*}}\right]=\infty$.
@@ Appendix E Other types of adjusters
In addition to the admissible adjusters described in Section 2.3, there exists another known class of admissible adjusters, originally proposed by Shafer et al., (2011): for $\mathfrak{e}\in[1,\infty]$,
@@ E.2 Generalized adjusters based on spines are not $\mathfrak{e}$-lifters
While Theorem 3 leaves open the possibility of other functional forms for adjusters, one generalization that does not work is the class of functions that take both the running maximum of a process and its current value. These functions have the form $\mathsf{A}:[1,\infty]\times[0,\infty]\to[0,\infty]$. Dawid et al., 2011a derive a characterization theorem for such generalized adjusters via univariate functions that they call spines. Based on spines, they give a family of admissible spine-based adjusters having the form
@@ Appendix F A bibliographic note on terminology
The notion of “lifting” in this paper is largely distinct from that of Getoor and Sharpe, (1972), who used the term in the (unrelated) context of continuous-time, complex-valued martingales. Their notion has since been called Hypothèse (Brémaud and Yor,, 1978), extension (Dubins et al.,, 1996), and immersion (Beghdadi-Sakrani and Emery,, 2006), among others, as it has been adapted to the (also unrelated) literature of Brownian innovations (e.g., Tsirelson,, 1998). Our terminology only concerns the lifting of events and evidence processes across filtrations defined in discrete time, with a focus on the validity of inference procedures at stopping times.
Further, as noted in Vovk, 2021a (, remark 1), it is a coincidence that the subject of Getoor and Sharpe, (1972)’s study was termed “conformal martingales,” which are unrelated to Vovk, 2021a ; Vovk, 2021b ’s “conformal test martingales” that we discuss in Example 1.
@@ G.1 Proof of Lemma 1
The direction b $\Rightarrow$ a is immediate. For the lifting direction a $\Rightarrow$ b, as alluded to in the main paper, the proof adapts the logic from Ramdas et al., (2020, lemma 1) and Howard et al., (2021, lemma 3) to stopping time validity across filtrations $\mathbb{F}$ and $\mathbb{G}$, where $\mathbb{G}\subseteq\mathbb{F}$.
@@ G.2 Proof of Theorem 2
First, since $(\mathfrak{e}_{t})_{t\geq 0}$ is an e-process for $\mathcal{P}$ in $\mathbb{G}$, we know from Ville’s inequality for e-processes (Ville,, 1939; Ramdas et al.,, 2020) that
@@
Next, given an adjuster $\mathsf{A}$, we can find its corresponding p-to-e calibrator (13), i.e., $\mathsf{C}(\mathfrak{p})=\mathsf{A}(1/\mathfrak{p})$ for $\mathfrak{p}\in[0,1]$. Then, for any $\mathbb{F}$-stopping time $\tau$, since $\mathfrak{p}_{\tau}^{*}$ is a valid p-value for $\mathcal{P}$, $\mathsf{C}(\mathfrak{p}_{\tau}^{*})$ is a valid e-value for $\mathcal{P}$ (see, e.g., Vovk and Wang,, 2021, proposition 2.1). In other words, for any $\mathbb{F}$-stopping time $\tau$,
@@ G.3 Two characterization lemmas and the proof of Theorem 3
($\Leftarrow$) Given any $P$, let $(M_{t}(P))_{t\geq 0}$ be a test supermartingale for $P$ in $\mathbb{G}$. By supermartingale optional stopping (e.g., Durrett,, 2019, theorem 4.8.4), $(M_{t}(P))_{t\geq 0}$ is an e-process for $\{P\}$ in $\mathbb{G}$. Then, condition (27) implies that there exists an e-process $(\mathfrak{e}_{t}^{\prime})_{t\geq 0}$ for $\{P\}$ in $\mathbb{G}$ satisfying $\mathsf{A}(M_{t}(P)^{*})\leq\mathfrak{e}_{t}^{\prime}$, $P$-a.s., for all $t\geq 0$. By Ramdas et al., (2020, lemma 6), there exists a test supermartingale $(M_{t}^{\prime}(P))_{t\geq 0}$ for $P$ in $\mathbb{G}$ that upper-bounds $(\mathfrak{e}_{t}^{\prime})_{t\geq 0}$, implying
@@
($\Rightarrow$) Suppose that $\mathsf{A}$ is an adjuster for test supermartingales, in the sense of (26). For any $\mathcal{P}$, consider an arbitrary e-process $(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in some filtration $\mathbb{G}$. By Ramdas et al., (2020, lemma 6, (i) $\Rightarrow$ (vi)), for each $P\in\mathcal{P}$, there is a test supermartingale $(M_{t}(P))_{t\geq 0}$ for $P$ in $\mathbb{G}$ that upper-bounds the e-process, that is,
@@
Finally, let $(\mathfrak{e}_{t}^{\prime})_{t\geq 0}$ be defined as $\mathfrak{e}_{0}^{\prime}=1$ and $\mathfrak{e}_{t}^{\prime}=\mathsf{A}(\mathfrak{e}_{t}^{*})$ for each $t\geq 1$. By Ramdas et al., (2020, lemma 6, (vi) $\Rightarrow$ (i)), equation (59) implies that $(\mathfrak{e}_{t}^{\prime})_{t\geq 0}$ is an e-process for $\mathcal{P}$ in $\mathbb{G}$, and it trivially satisfies condition (27). ∎
@@
The equivalence between a and b follows from Dawid et al., 2011b (, theorem 1). Next, Theorem 2 establishes that a implies e, and e trivially implies d once we set $\mathbb{F}=\mathbb{G}$. Then, the “if” direction of Lemma 3 shows that d implies 27. Finally, the “if” direction of Lemma 2 establishes that 27 $\Rightarrow$ b. Put together, we showed that b $\Leftrightarrow$ a $\Rightarrow$ e $\Rightarrow$ d $\Rightarrow$ 27 $\Rightarrow$ b, proving that all statements are equivalent. ∎
@@ H.1 How much power do we actually lose?
We can also plot both the statistical power of the induced test, as well as the e-power (Vovk and Wang,, 2024; Ramdas and Wang,, 2025), against the sample size. Given an alternative distribution $Q$, we define the e-power of an e-process $(\mathfrak{e}_{t})_{t\geq 0}$ at time $t$ is defined as $\mathbb{E}_{Q}[\log\mathfrak{e}_{t}]$. It provides a different notion of statistical power focusing on how quickly the e-process can grow under the alternative.
@@ H.2 An empirical comparison with existing approaches for the $h$-step-ahead forecast evaluation problem
Due to the unique setup of the problem—specifically, due to Arnold et al., (2023)’s filtration construction—if we only need a sequential test then there exist alternative approaches that bypass the need to construct an e-process (and thus an adjuster). Arnold et al., (2023)’s strategy is the following: given $h$ e-processes $\mathfrak{e}^{[k]}$ for $\mathcal{P}$ in a sub-filtration $\mathbb{F}^{[k]}\subsetneq\mathbb{F}$, the unadjusted mean of the e-processes $\mathfrak{e}^{[k]}$, call it the “simple mean” $\bar{M}_{t}=\frac{1}{h}\sum_{k=1}^{h}\mathfrak{e}_{t}^{[k]}$ (not an e-process), satisfies a lagged version of $\mathbb{F}$-anytime-validity in the form of
@@
To obtain a fully anytime-valid test, prior works also note that one can utilize the harmonic-mean p-merging function (Vovk and Wang,, 2020) in the $\mathfrak{p}$-domain:
@@
To construct an e-process, the only other known way (Choe and Ramdas,, 2023, ppn. 5.3) is to calibrate the harmonic-mean p-process in the p-to-e direction: use
@@
for some p-to-e calibrator $\mathsf{C}$. Yet, this approach is suboptimal. At a high level, this is because the harmonic-mean p-merging function, like any other admissible p-merging function, implicitly transforms each p-value into an e-value using different p-to-e calibrators before averaging them (Gasparin et al.,, 2024). This implies that Choe and Ramdas, (2023)’s approach (64) effectively applies p-to-e calibration twice. In contrast, the adjust-then-combine approach (25) only goes through one p-to-e calibrator (by the use of an adjuster).

### v5

@@ D.1 Illustrative case: Sequentially testing a bounded IID mean
For this test martingale, it is possible to derive the conditions with which the running maximum of the test martingale $(M_{t})_{t\geq 0}$ tends to infinity as $t\to\infty$, even under the null. Let $M_{t}^{*}=\sup_{i\leq t}M_{i}$ denote the running maximum of the test martingale. The result below follows from an anti-concentration lemma due to Ramdas et al. (2020).
Proposition 1 (Divergence of the expected running maximum for test martingales).
Suppose that the $\mathbb{F}$-predictable sequence $(\lambda_{t})_{t\geq 0}$ is chosen such that, under each $P\in\mathcal{P}^{\mu}$, the sum of quadratic variations in the data, weighted by $\lambda_{t}^{2}$, tends to infinity:
@@
A practically relevant case is when the sequence decays at a rate of $\left({\sqrt{t\log t}}\right)^{-1}$ (or a slower rate). For example, Waudby-Smith and Ramdas (2024) recommend using $\hat{\lambda}_{t}\propto\left(\sqrt{\hat{\sigma}_{t-1}^{2}t\log(t+1)}\right)^{-1}$, where $\hat{\sigma}_{t-1}^{2}$ is a $\mathcal{F}_{t-1}$-measurable, consistent estimate of the variance $\sigma^{2}$. In such cases, we may use classical generalizations of the SLLN to martingales involving predictable weights (see, e.g., theorem 2.18 of Hall and Heyde, 1980) to show that condition (41) is met.
@@
We also know, by construction, each $Y_{t}$ is bounded by the constant $1+\frac{1-\mu}{\mu}$ (note that $\mu$ is fixed for each testing problem). Writing $\epsilon=\frac{1-\mu}{\mu}>0$, we have
@@
These two conditions, (44) and (45), are sufficient for an “anti-concentration” lemma for test martingales (Ramdas et al., 2020, lemma 33). In particular, the conditions imply that
$P\left(\sup_{t}M_{t}\geq x\right)\geq\frac{1}{(1+\epsilon)x},\quad\forall x\in(1,\infty).$ (46)
This is enough to show that the running maximum process $M_{t}^{*}=\sup_{i\leq t}M_{i}$ tends to infinity in expectation. More precisely, given a fixed $\mu\in(0,1)$,
$\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}^{*}}\right]\geq\mathbb{E}_{{P}}\left[{\lim_{t\to\infty}M_{t}^{*}}\right]=\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]=\int_{1}^{\infty}P\left(\sup_{t}M_{t}\geq x\right)dx=\int_{1}^{\infty}\frac{dx}{(1+\epsilon)x}=\infty,$ (47)
@@ D.2 When does this happen for test martingales (and e-processes)?
This phenomenon is not specific to the bounded IID mean testing problem—in fact, we may apply the anti-concentration lemma to any test martingale whose sum of quadratic variations is infinite. Many test martingales used in practice will satisfy this requirement: game-theoretically, as long as one does not decrease one’s bets “too quickly” (or stop betting entirely), which is the case for many powerful test martingales and e-processes, the variance in data will force the sum to infinity. Note that, for test martingales having the form (40), the bets correspond to the predictable sequence $(\lambda_{t})_{t\geq 1}$. More generally for e-processes, we remark that e-processes for composite nulls $\mathcal{P}$ are often constructed as the minimum of powerful test martingales for each null distribution $P\in\mathcal{P}$ (e.g., the UI e-process of Wasserman et al., 2020).
@@
By the first half of the anti-concentration lemma (Ramdas et al., 2020, lemma 33), we know
@@
Now, suppose for contradiction that $\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]$ were finite. Denoting $M_{\infty}^{*}=\lim_{t\to\infty}M_{t}^{*}=\sup_{t}M_{t}$, we see that $(M_{t})_{t\geq 0}$ is dominated by $M_{\infty}^{*}\in L^{1}$, and thus it is uniformly integrable (see, e.g., theorem 4.6.7 of Durrett, 2019). This then implies that the martingale $(M_{t})_{t\geq 0}$ converges to $M_{\infty}$ in $L^{1}$. In particular, $\mathbb{E}_{{P}}\left[{M_{\infty}}\right]=\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}}\right]=\mathbb{E}_{{P}}\left[{M_{0}}\right]=1$, which contradicts equation (49). Therefore, $\mathbb{E}_{{P}}\left[{\sup_{t}M_{t}}\right]$ must be infinite, and by Fatou’s lemma, $\lim_{t\to\infty}\mathbb{E}_{{P}}\left[{M_{t}^{*}}\right]=\infty$.
@@ Appendix E Other types of adjusters
In addition to the admissible adjusters described in Section 2.3, there exists another known class of admissible adjusters, originally proposed by Shafer et al. (2011): for $\mathfrak{e}\in[1,\infty]$,
@@ E.2 Generalized adjusters based on spines are not $\mathfrak{e}$-lifters
While Theorem 3 leaves open the possibility of other functional forms for adjusters, one generalization that does not work is the class of functions that take both the running maximum of a process and its current value. These functions have the form $\mathsf{A}:[1,\infty]\times[0,\infty]\to[0,\infty]$. Dawid et al. (2011a) derive a characterization theorem for such generalized adjusters via univariate functions that they call spines. Based on spines, they give a family of admissible spine-based adjusters having the form
@@ Appendix F A bibliographic note on terminology
The notion of “lifting” in this paper is largely distinct from that of Getoor and Sharpe (1972), who used the term in the (unrelated) context of continuous-time, complex-valued martingales. Their notion has since been called Hypothèse (Brémaud and Yor, 1978), extension (Dubins et al., 1996), and immersion (Beghdadi-Sakrani and Emery, 2006), among others, as it has been adapted to the (also unrelated) literature of Brownian innovations (e.g., Tsirelson, 1998). Our terminology only concerns the lifting of events and evidence processes across filtrations defined in discrete time, with a focus on the validity of inference procedures at stopping times.
Further, as noted in Vovk (2021a, remark 1), it is a coincidence that the subject of Getoor and Sharpe (1972)’s study was termed “conformal martingales,” which are unrelated to Vovk (2021a, b)’s “conformal test martingales” that we discuss in Example 1.
@@ G.1 Proof of Lemma 1
The direction b $\Rightarrow$ a is immediate. For the lifting direction a $\Rightarrow$ b, as alluded to in the main paper, the proof adapts the logic from Ramdas et al. (2020, lemma 1) and Howard et al. (2021, lemma 3) to stopping time validity across filtrations $\mathbb{F}$ and $\mathbb{G}$, where $\mathbb{G}\subseteq\mathbb{F}$.
@@ G.2 Proof of Theorem 2
First, since $(\mathfrak{e}_{t})_{t\geq 0}$ is an e-process for $\mathcal{P}$ in $\mathbb{G}$, we know from Ville’s inequality for e-processes (Ville, 1939; Ramdas et al., 2020) that
@@
Next, given an adjuster $\mathsf{A}$, we can find its corresponding p-to-e calibrator (13), i.e., $\mathsf{C}(\mathfrak{p})=\mathsf{A}(1/\mathfrak{p})$ for $\mathfrak{p}\in[0,1]$. Then, for any $\mathbb{F}$-stopping time $\tau$, since $\mathfrak{p}_{\tau}^{*}$ is a valid p-value for $\mathcal{P}$, $\mathsf{C}(\mathfrak{p}_{\tau}^{*})$ is a valid e-value for $\mathcal{P}$ (see, e.g., Vovk and Wang, 2021, proposition 2.1). In other words, for any $\mathbb{F}$-stopping time $\tau$,
@@ G.3 Two characterization lemmas and the proof of Theorem 3
($\Leftarrow$) Given any $P$, let $(M_{t}(P))_{t\geq 0}$ be a test supermartingale for $P$ in $\mathbb{G}$. By supermartingale optional stopping (e.g., Durrett, 2019, theorem 4.8.4), $(M_{t}(P))_{t\geq 0}$ is an e-process for $\{P\}$ in $\mathbb{G}$. Then, condition (27) implies that there exists an e-process $(\mathfrak{e}_{t}^{\prime})_{t\geq 0}$ for $\{P\}$ in $\mathbb{G}$ satisfying $\mathsf{A}(M_{t}(P)^{*})\leq\mathfrak{e}_{t}^{\prime}$, $P$-a.s., for all $t\geq 0$. By Ramdas et al. (2020, lemma 6), there exists a test supermartingale $(M_{t}^{\prime}(P))_{t\geq 0}$ for $P$ in $\mathbb{G}$ that upper-bounds $(\mathfrak{e}_{t}^{\prime})_{t\geq 0}$, implying
@@
($\Rightarrow$) Suppose that $\mathsf{A}$ is an adjuster for test supermartingales, in the sense of (26). For any $\mathcal{P}$, consider an arbitrary e-process $(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in some filtration $\mathbb{G}$. By Ramdas et al. (2020, lemma 6, (i) $\Rightarrow$ (vi)), for each $P\in\mathcal{P}$, there is a test supermartingale $(M_{t}(P))_{t\geq 0}$ for $P$ in $\mathbb{G}$ that upper-bounds the e-process, that is,
@@
Finally, let $(\mathfrak{e}_{t}^{\prime})_{t\geq 0}$ be defined as $\mathfrak{e}_{0}^{\prime}=1$ and $\mathfrak{e}_{t}^{\prime}=\mathsf{A}(\mathfrak{e}_{t}^{*})$ for each $t\geq 1$. By Ramdas et al. (2020, lemma 6, (vi) $\Rightarrow$ (i)), equation (59) implies that $(\mathfrak{e}_{t}^{\prime})_{t\geq 0}$ is an e-process for $\mathcal{P}$ in $\mathbb{G}$, and it trivially satisfies condition (27). ∎
@@
The equivalence between a and b follows from Dawid et al. (2011b, theorem 1). Next, Theorem 2 establishes that a implies e, and e trivially implies d once we set $\mathbb{F}=\mathbb{G}$. Then, the “if” direction of Lemma 3 shows that d implies 27. Finally, the “if” direction of Lemma 2 establishes that 27 $\Rightarrow$ b. Put together, we showed that b $\Leftrightarrow$ a $\Rightarrow$ e $\Rightarrow$ d $\Rightarrow$ 27 $\Rightarrow$ b, proving that all statements are equivalent. ∎
@@ H.1 How much power do we actually lose?
We can also plot the power of the induced test, as well as the e-power (Vovk and Wang, 2024; Ramdas and Wang, 2025), against the sample size. Given an alternative distribution $Q$, we define the e-power of an e-process $(\mathfrak{e}_{t})_{t\geq 0}$ at time $t$ is defined as $\mathbb{E}_{Q}[\log\mathfrak{e}_{t}]$. It provides a different notion of statistical power focusing on how quickly the e-process can grow under the alternative.
@@ H.2 An empirical comparison with existing approaches for the $h$-step-ahead forecast evaluation problem
Due to the unique setup of the problem—specifically, due to Arnold et al. (2023)’s filtration construction—if we only need a sequential test then there exist alternative approaches that bypass the need to construct an e-process (and thus an adjuster). Arnold et al. (2023)’s strategy is the following: given $h$ e-processes $\mathfrak{e}^{[k]}$ for $\mathcal{P}$ in a sub-filtration $\mathbb{F}^{[k]}\subsetneq\mathbb{F}$, the unadjusted mean of the e-processes $\mathfrak{e}^{[k]}$, call it the “simple mean” $\bar{M}_{t}=\frac{1}{h}\sum_{k=1}^{h}\mathfrak{e}_{t}^{[k]}$ (not an e-process), satisfies a lagged version of $\mathbb{F}$-anytime-validity in the form of
@@
To obtain a fully anytime-valid test, prior works also note that one can utilize the harmonic-mean p-merging function (Vovk and Wang, 2020) in the $\mathfrak{p}$-domain:
@@
To construct an e-process, the only other known way (Choe and Ramdas, 2023, ppn. 5.3) is to calibrate the harmonic-mean p-process in the p-to-e direction: use
@@
for some p-to-e calibrator $\mathsf{C}$. Yet, this approach is suboptimal. At a high level, this is because the harmonic-mean p-merging function, like any other admissible p-merging function, implicitly transforms each p-value into an e-value using different p-to-e calibrators before averaging them (Gasparin et al., 2025). This implies that Choe and Ramdas (2023)’s approach (64) effectively applies p-to-e calibration twice. In contrast, the adjust-then-combine approach (25) only goes through one p-to-e calibrator (by the use of an adjuster).